the distance between the lines3x+4y=9and 6x+8y+15=0 is??

Grade:12th pass

1 Answers

Vikas TU
7 years ago
distance between two parallel lines is given by the formulae:
m = |9 – 15|/|(root(36 + 64)| = 6/10 = 3/5 is the answer.